OpenStudy (sparklyme):

Use the quadratic formula to find the x-intercepts: y=-(x-10)^2+140

OpenStudy (welshfella):

first rearrange the equaton to the form ax^2 + bx + c = 0

OpenStudy (mathstudent55):

The x-intercepts occur where y = 0.

OpenStudy (mathstudent55):

First, replace y with 0. Then write the quadratic in standard form, as @welshfella showed above. Then you can read the coefficients a, b, and c, and use the quadratic formula.

OpenStudy (sparklyme):

@welshfella @mathstudent55 So would it be\[-x ^{2}-100+140=0\]

OpenStudy (welshfella):

-(x - 10)^2 + 140 = - ( x^2 - 20x + 100) + 140 = -x^2 + 20x - 100 + 140 = -x^2 + 20x + 40

OpenStudy (sparklyme):

@welshfella @mathstudent55 where did you get the 20x?

OpenStudy (mathstudent55):

Here it is step by step: \(y=-(x-10)^2+140\) \(0 = -(x-10)^2+140\) \(-(x-10)^2+140 =0\) \(-(x^2 - 20x + 100) + 140 = 0\) \(-x^2 + 20x - 100 + 140 = 0\) \(-x^2 + 20x + 40 = 0\) \(x^2 - 20x - 40 = 0\)

OpenStudy (sparklyme):

Where did you get 20x?

OpenStudy (mathstudent55):

Notice how I dealt with \((x - 10)^2\) It is the square of a binomial and follows the pattern: \((a - b)^2 = a^2 - 2ab + b^2\)

OpenStudy (mathstudent55):

If you don't remember that, you can always go back to \((a - b)^2 = (a - b)(a - b)\) and use FOIL

OpenStudy (sparklyme):

Yea. I like that better thanks.

OpenStudy (sparklyme):

So what would I FOIL? The whole thing or just what is inside the ()?

OpenStudy (mathstudent55):

You just use a product of binomials to handle the square of the binomial. \((x - 10)(x - 10) = x^2 -10x - 10x + 100 = x^2 - 20x + 100\)

OpenStudy (sparklyme):

ohhh....right. thank you so much for everything. it means a lot!

OpenStudy (mathstudent55):

I'll show you the steps again using the product of the binomials instead of the square of the binomial.

OpenStudy (sparklyme):

Okay!

OpenStudy (mathstudent55):

We start out just as we did before. \(y=-(x-10)^2+140\) Replace y with 0. \(0 = -(x-10)^2+140\) Switch sides to have zero on the right side. \(-(x-10)^2+140 = 0\) Now show the product of binomials instead of the square of the binomial. \(- (x - 10)(x - 10) +140 = 0\)

OpenStudy (mathstudent55):

Now we use FOIL. Remember we must keep that negative sign to the left of the the product of the binomials. We will deal with the negative sign after we use FOIL. Here is the last line from above again. \(- \color{red}{(x - 10)(x - 10)} +140 = 0\) Now we use FOIL keeping the negative sign to the left as it must be kept. \(- \color{red}{(x^2 - 10x -10x + 100)} +140 = 0\) Now we combine like terms after FOIL. \(- \color{red}{(x^2 - 20x + 100)} +140 = 0\) Now we deal with the negative outside the parentheses. To distribute a negative sign, think of the negative sign as being -1 multiplying the parentheses. We multiply everything inside the parentheses by -1. What that does is change every sign inside the parentheses. \(- x^2 + 20x - 100 +140 = 0\) Ok so far?

OpenStudy (sparklyme):

Yes.

OpenStudy (sparklyme):

After this I think I know what to do. Thank you so much for everything!

OpenStudy (mathstudent55):

Now we combine like terms, -100 + 140 = 40 \(- x^2 + 20x +40 = 0\) Now I multiply everything on both sides by -1 to get rid of the negative sign on the x^2 term. \(x^2 - 20x - 40 = 0\) This last step is just my preference. You can use the previous line with first term -x^2 already in the quadratic formula. Now you are ready for the quadratic formula.

OpenStudy (mathstudent55):

You're welcome.